# Model predictive control is almost optimal for restless bandits

### v1

@@ 6 Numerical illustrations
In this section, we illustrate numerically the performance of the LP-update policy. We show that it performs very well in practice and outperform classical heuristics in most cases. We choose to compare against two heuristics of the literature: the LP-index policy of Gast et al. (2023a) and the FTVA of Hong et al. (2023). We choose these two heuristics because they are natural and simple to implement and do not rely on hard-to-tune hyperparameters (further discussion for our choice can be found in Appendix  F). All parameters of the examples are provided in Appendix F. All codes are provided in supplementary material. In the future, a Github link will be provided.
@@
For our first comparison, we consider three representative examples: a randomly generated example (in dimension 8), the main example used in Hong et al. (2023) and described in their Appendix G.2 (abbreviated Example Hong et al. (2023) in the following), and Example 2 of Figure 7.4 of Chen (2022) (abbreviated Example Chen (2022)). We report the result in Figure 1. We observe that in all three cases, both the LP-update and the FTVA policies are asymptotically optimal, but the LP-update policy outperforms FTVA substantially. The situation for LP-priority is quite different: as shown in Gast et al. (2023a), in dimension 8, the LP-priority policy is asymptotically optimal for most of the randomly generated examples. This is the case for example (a) of Figure 1 for which LP-update and LP-priority give an essentially equivalent performance and are essentially almost optimal.
Note that the authors of Chen (2022) provide the numerical value of the optimal policy $N$ for Example Chen (2022), which can be done because $|\mathcal{S}|=3$. We observe that the LP-update is very close to this value.
Refer to caption Refer to caption Refer to caption
(a) Random example (b) Example Hong et al. (2023) (c) Example Chen (2022)
Figure 1: Performance as a function of $N$
@@
To explore the difference of performance between the LP-update policy and FTVA, we study the dynamics as a function of time of the different policies for the first two examples of Figure 1. We plot in Figure 2 the evolution of the system as a function of time (for a single trajectory). We observe that if the distance between $X_{t}$ and $x^{*}$ are similar for LP-update and FTVA, the rotated cost is much smaller for LP-update, which explains its good performance.
@@
To explore the system dynamics in more details, we focus on Example Chen (2022) and study the behavior of the LP-update, FTVA and LP-priority policy. In Figures 3(a,b,c), we present a trajectory of each of the three policies: each orange point corresponds to a value of $X_{t}\in\Delta_{\mathcal{S}}$ (as this example is in dimension $|\mathcal{S}|=3$, the simplex $\Delta_{\mathcal{S}}$ can be represented as a triangle). This example has an unstable fixed point (Assumption 3 is not satisfied). Both LP-update and FTVA concentrate their behavior around this fixed point but the LP-priority policy exhibits two modes. When concentrating on the rotated cost (Figure 3(d)), we observe that it is much smaller for LP-update when compared to FTVA. This explains why LP-update performs better as shown in Figure 1(c).
@@
In Figure 4, we study the influence of different parameters on the performance of LP-update. In each case, we generated $20$ examples and plot the average “normalized” performance among these $20$ examples. By “normalized”, we mean that for each example, we divide the performance of the policy by the value of the LP problem (5). The first plot 4(a) shows that the influence of the parameter $\tau$ is marginal (the curves are not distinguishable for $\tau\in\{3,5,10\}$). In the second plot 4(b), we observe that the sub-optimality gap of LP-update does not seem to depend much on the state space size whereas the one of FTVA does: FTVA performs worse when $|\mathcal{S}|$ is large probably because there are fewer synchronized arms. In the last plot 4(c), we study the performance as a function of the budget $\alpha$.
@@ 7 Conclusion
In this paper, we study the problem of constructing an efficient policy for restless multi-armed bandit for the average reward case. We show that under quite general condition, a simple model-predictive control algorithm provides a solution that achieves both the best known upper bound up to know ($O(1/\sqrt{N})$ or $\exp(-\Omega{N})$ for stable non-degenerate problems), but also works very efficiently in practice. Our paper provides the first analysis of this policy for the average reward criterion.
Our paper uses a novel framework based on dissipativity that helps up to make a crisp connection between the finite- and the infinite-horizon problems and we are the first to us it in this context. We believe that this framework is what makes our analysis simple and easily generalizable to other settings. As an example, we discuss in Appendix B.2 potential generalization to the multi-action multi-constraint bandits, that is an interesting direction for future work.
@@ Appendix A Additional Literature Review
In his seminal work Whittle (1988) generalized the problem to the restless multi-armed bandit problem in its modern form and conjectured that under a condition known as indexability, an index policy is asymptotically optimal (in the number of arms) for the RMAB problem. Weber & Weiss (1990) showed a counter example to this conjecture, further going on to show in the same work that the conjecture did hold under a global attractor and ergodicity condition of a single-arm. This result was generalized to the multi-action setting by Hodge & Glazebrook (2015). On the other hand, Verloop (2016); Gast et al. (2023b) generalized the set of index policies by constructing a set of priority policies known as LP priority index. More surprisingly, Gast et al. (2023a; b) showed these priority policies are in fact exponentially close to being optimal, under three conditions: uniform global attractor property (UGAP), non-degeneracy of the fixed point and global exponential stability Gast et al. (2023a). The restriction of indexability is removed in Gast et al. (2023b) and the condition of UGAP was first removed Hong et al. (2024a).
Gast et al. (2023a; b) observed that under the conditions of non-degeneracy, there exists a local neighborhood around the fixed point where the priority policy set is affine. This is the key observation which leads to the exponentially small error bounds under simple local stability conditions. This observation around the fixed point drove several works (Hong et al., 2023; 2024a; 2024b; Goldsztajn & Avrachenkov, 2024; Yan, 2024) to shift their perspective towards policies that look to steer the dynamical system towards the optimal fixed point. Hong et al. (2023) created a virtual system that was driven to the fixed point. To transform this into an asymptotically optimal policy, they required the real system to synchronize with the virtual system, hence, they developed the synchronization assumption. Unlike Hong et al. (2023) which passively allowed arms to align with each other eventually, Hong et al. (2024a) used a two set policy to actively align non aligned arms to the fixed point. In contrast to this form of control, Yan (2024) designed the align and steer policy to steer the mean field dynamical control system towards the fixed point. Yan (2024)’s policy is asymptotically optimal under the assumptions of controllability of the dynamical system to more directly move the state of the system towards the fixed point over time. In parallel to this work, Goldsztajn & Avrachenkov (2024) used a similar idea of mean field dynamical control for weakly coupled systems under a relaxed but generalized constraint set if the single arm process is unichain and aperiodic.
@@
There are also a lot of recent papers on RMABs for the finite-horzon reward criteria Hu & Frazier (2017); Zayas-Caban et al. (2019); Brown & Smith (2020); Ghosh et al. (2022); Zhang & Frazier (2021); Gast et al. (2023a), or the infinite-horizon discounted reward Zhang & Frazier (2022); Ghosh et al. (2022). Our results are close in spirit to those since our policy shows how to use a finite-horizon policy to construct a policy that is asymptotically optimal for the infinite-horizon rewards.
@@ B.1 Rounding procedure
    First, given $\bm{u}=\mu_{\tau}({\bm{x}}(\textbf{S}(t)))$, we need to construct a vector $\bm{U}^{N}$ that is as close as possible to $\bm{u}$ while satisfying the constraints that $\sum_{i}u_{i}<=\alpha$ and $N\bm{U}^{N}$ is an integer.
@@
This leads us to construct Algorithm 2. This algorithm first construct a $\bm{v}$ that is as close as possible to $\bm{u}$ while satisfying $\|\bm{v}\|\leq\alpha$. There might be multiple choices for this first step and the algorithm can choose any. For instance, applied to the example such that ${\bm{x}}=(10,10,10,9)$ and $\bm{u}=(10,4.9,4.6,0)$, this algorithm would first construct a vector $\bm{v}$ that can be any convex combination of $(10,4.4,4.6,0,0)$ and $(10,4.9,4.1,0,0)$. For instance, if the algorithm chose the vector $\bm{v}=(10,4.4,4.6,0)$, then the algorithm would produce $(10,5,4,0)$ with probability $0.4$ and $(10,4,5,0)$ with probability $0.6$. Once this is done, the algorithm outputs a $\bm{U}^{N}$ that satisfies the constraints and such that $\mathbb{E}[\bm{U}^{N}]=\bm{v}$. In some cases, there might be multiple distributions that work. The algorithm may output any. An efficient procedure to implemented this last step is provided in Section 5.2.3 of Ioannidis & Yeh (2016).
@@ C.1 Part 1, Properties of the dynamical control problem 8:
Proof.
Now note, $\textbf{y}\cdot\textbf{1}=1$, where 1 is the all $1^{\prime}$s vector. We will overload the notation slightly by letting $\textbf{y}(t)$ be the optimal trajectory taken by the optimal policy $\pi^{*}_{\beta}$ for 20. Hence, for any constant $g$ we have:
@@
More importantly, by plugging these definitions into (29) and making the dependence of y on the maximal policy, we obtain the following recursion equation:
@@
which completes the proof. ∎
@@ C.2 Part 2: Dissipativity
which implies the results of the lemma. ∎
@@ C.3 Part 3: MPC is optimal for the deterministic problem
For all ${\bm{x}}\in\Delta_{\mathcal{S}}$, for any $\epsilon>0$, there exists $t$ such that:
$L_{t}({\bm{x}})-L_{t-1}(\Phi({\bm{x}},\bm{u}))<\epsilon.$
where $\bm{u}\in\mathcal{U}({\bm{x}})$ is $\mu_{t}({\bm{x}})$.
Proof.
@@
$\sum_{t=0}^{T-1}\tilde{l}({\bm{x}}(t),\bm{u}(t))$ For $L_{\tau}$
$\sum_{t=0}^{T-1}R({\bm{x}}(t),\bm{u}(t))+\lambda({\bm{x}}(\tau)).$ For $W_{\tau}$
By Lemma  C.3, the rotated cost is $\tilde{l}(\textbf{x},\bm{u})=g^{*}-R({\bm{x}},\bm{u})+\lambda({\bm{x}})-\lambda(\Phi({\bm{x}},\bm{u}))$. As any valid control of $L_{\tau}$ and $W_{\tau}$ satisfy ${\bm{x}}(t+1)=\Phi({\bm{x}}(t),\bm{u}(t))$, the objective of $L_{\tau}$ can be rewritten as:
@@
As the last two terms correspond to the objective function for $W_{\tau}$, this shows that $L_{\tau}({\bm{x}})=\tau g^{\star}+\lambda\cdot{\bm{x}}-W_{\tau}{({\bm{x}})}$. Clearly, the two objectives are equivalent and $\mu_{\tau}({\bm{x}})=\bm{u}(0)$. Now note, by Lemma  C.2 we have,
@@
is bounded. Due to dissipativity, $L_{\tau}(\textbf{x})$ is monotone increasing in $\tau$. It follows that
$L_{\tau}({\bm{x}})-L_{\tau-1}(\Phi({\bm{x}},\bm{u}))<\epsilon.$
This completes the proof. ∎
C.4 Proof of Theorem 4.1: Computation details
@@
$\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[L_{\tau}({\bm{X}(t)})-L_{\tau}(\Phi({\bm{X}(t)},\bm{u}(t)))+\epsilon-\lambda\cdot{\bm{X}}(t)+\lambda\cdot\Phi({\bm{X}}(t),\bm{u}(t))\right],$ (35)
where we use the definition of the rotated cost for the first equality, the dynamic principle for the second line and the identity (17) for the last time.
$\frac{1}{T}\sum_{t=0}^{T-1}[L_{\tau}({\bm{X}(t)})-\lambda\cdot{\bm{X}(t)}]$ $=\frac{1}{T}\sum_{t=0}^{T-1}[L_{\tau}({\bm{X}(t+1)})-\lambda\cdot{\bm{X}(t+1)}]$
$\quad+\frac{1}{T}\left(L_{\tau}({\bm{X}(0)})-L_{\tau}({\bm{X}(T)})+\lambda\cdot{\bm{X}(0)}-\lambda\cdot{\bm{X}(T)}\right)$
@@
$\epsilon+\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\left[L_{\tau}({\bm{X}}(t+1))-L_{\tau}(\Phi({\bm{X}}(t),\bm{u}(t)))-\lambda\cdot{\bm{X}}(t+1)+\lambda\cdot\Phi({\bm{X}}(t),\bm{u}(t))\right],$
which shows (18)
@@
The lemma can then be proven by adapting the proof of Theorem 3 of Gast et al. (2023b) and in particular their Equation (17). This is proven by using Lemma 1 Gast et al. (2023b) that implies that $\Pr[\|{\bm{X}(t+1)}-\Phi({\bm{X}(t)},\bm{U}(t))\|\geq\epsilon]\leq e^{-C^{\prime}N}$. ∎
We are now ready to prove Theorem 4.2. We can follow the proof of Theorem 4.1 up to (18) that shows that $V_{\text{opt}}^{N}(\textbf{x})-V_{\text{LP},\tau}^{(N)}(\textbf{x})$ is bounded [^4] In this expression, the term $(\alpha N-\lfloor{\alpha N}\rfloor)/N$ is equal to 00 here because we assumed that $\alpha N$ is an integer. by
@@
Proof Outline
This section is dedicated to proving the Lipschitz property for the value function $V^{\infty}({\bm{x}})$ in ${\bm{x}}$ under the ergodicity Assumption 1. For simplicity of exposition we will set $k=1$, although it is not hard to extrapolate the proof for $k>1$ and is left as an exercise for the reader.

### v2

@@ 6 Numerical illustrations
We illustrate numerically the performance of the LP-update policy, showing that it performs very well in practice and outperforms classical heuristics in most cases. We choose to compare against two heuristics of the literature: the LP-priority policy of Gast et al. (2023a) and the follow the virtual advice policy (FTVA) of Hong et al. (2023). The LP-priority policy computes a solution to (5) and uses it to compute a priority order on states, arms are then pulled according to this priority order. The LP-priority policy is an index policy —similarly to Whittle index policy— but has the advantage of being always defined, whereas Whittle index policy is only defined under a technical condition known as indexability. The LP-update policy has only been shown to be optimal under the UGAP assumption. FTVA constructs a virtual collection of $N$ arms which follow the solution to (5). When a virtual arm is pulled, the corresponding real arm is pulled if the budget allows it, otherwise the real arm is left alone. If the virtual and real arm choose the same actions, they evolve identically, otherwise independently. We choose these two heuristics because they are natural and simple to implement and do not rely on hard-to-tune hyperparameters (further discussion for our choice can be found in the supplementary material along with choices for the parameters, see Appendix  G). We provide the parameters in Appendix G.All codes are provided in supplementary material.
@@
For our first comparison, we consider three representative examples: a randomly generated example (in dimension 8), the main example used in Hong et al. (2023) and described in their Appendix G.2 (abbreviated Example Hong et al. (2023) in the following), and Example 2 of Figure 7.4 of Yan (2022) (abbreviated Example Yan (2022)). We report the result in Figure 1. We observe that in all three cases, both the LP-update and the FTVA policies are asymptotically optimal, but the LP-update policy outperforms FTVA substantially. The situation for LP-priority is quite different: as shown in Gast et al. (2023a), in dimension 8, the LP-priority policy is asymptotically optimal for most of the randomly generated examples. This is the case for example (a) of Figure 1 for which LP-update and LP-priority give an essentially equivalent performance and are essentially almost optimal. The authors of Yan (2022) provide the numerical value of the optimal policy for Example Yan (2022), which can be done because $|\mathcal{S}|=3$. The LP-update is very close to this value.
@@
To explore the difference of performance between the LP-update policy and FTVA, we study the dynamics of the different policies for the first two examples of Figure 1. Figure 2 plots the evolution of the system over time (for a single trajectory). We observe that if the distance between $X_{t}$ and $x^{*}$ are similar for LP-update and FTVA, the rotated cost is much smaller for LP-update, which explains its good performance.
@@
To explore the system dynamics in more details, we focus on Example Yan (2022) and study the behavior of the LP-update, FTVA and LP-priority policy. In Figures 3(a,b,c), we present a trajectory of each of the three policies: each orange point corresponds to a value of $X_{t}\in\Delta_{\mathcal{S}}$ (as this example is in dimension $|\mathcal{S}|=3$, the simplex $\Delta_{\mathcal{S}}$ can be represented as a triangle). This example has an unstable fixed point (Assumption 3 is not satisfied). Both LP-update and FTVA concentrate their behavior around this fixed point but the LP-priority policy exhibits two modes. The rotated cost (Figure 3(d)) is much smaller for LP-update when compared to FTVA. This explains why the LP-update performs better as shown in Figure 1(c).
@@
Figure 4, studies the influence of different parameters on the performance of LP-update. In each case, we generated $20$ examples and plot the average “normalized” performance among these examples. “normalized” here means we divide the performance of the policy by the value of the LP problem (5). The first plot 4(a) shows that the influence of the parameter $\tau$ is marginal (the curves are not distinguishable for $\tau\in\{3,5,10\}$). Plot 4(b) indicates that the sub-optimality gap of LP-update is not too sensitive to the state space size unlike FTVA which degrades when $|\mathcal{S}|$ is large probably because there are fewer synchronized arms. The last plot 4(c) studies the performance as a function of the budget parameter, $\alpha$.
@@ 7 Conclusion
In this paper, we study the problem of constructing an efficient policy for restless multi-armed bandit for the average reward case. We show that under a quite general condition, a simple model-predictive control algorithm provides a solution that achieves both the best known upper bound ($O(1/\sqrt{N})$ or $\exp(-\Omega(N))$ for stable non-degenerate problems), and also works very efficiently in practice. Our paper provides the first analysis of this policy for the average reward criterion. By using a novel framework based on dissipativity we are able to make a subtle connection between the finite and the infinite-horizon problems and we are the first to use it in this context. We believe that this framework is what makes our analysis simple and easily generalizable to other settings. As an example, we discuss in Appendix B.2 the generalization to multi-action multi-constraint bandits. This connection also lends itself to a model-based learning setting. Furthermore, we believe that this framework could be used in the heterogeneous bandit problem. These problems are a matter for future work.
@@ Appendix A Additional Literature Review
In his seminal work Whittle (1988) generalized the problem to the restless multi-armed bandit problem in its modern form and conjectured that under a condition known as indexability, an index policy is asymptotically optimal (in the number of arms) for the RMAB problem. Weber and Weiss (1990) showed a counter example to this conjecture, further going on to show in the same work that the conjecture did hold under a global attractor (UGAP assumption) and ergodicity condition of a single-arm. This result was generalized to the multi-action setting by Hodge and Glazebrook (2015). On the other hand, Verloop (2016); Gast et al. (2023b) generalized the set of index policies by constructing a set of priority policies known as LP priority index while maintaining the UGAP assumption. More surprisingly, Gast et al. (2023a, b) showed these priority policies are in fact exponentially close to being optimal, under three conditions: uniform global attractor property (UGAP), non-degeneracy of the fixed point and global exponential stability Gast et al. (2023a). The restriction of indexability is removed in Gast et al. (2023b) and the condition of UGAP was first removed Hong et al. (2024a).
Gast et al. (2023a, b) observed that under the conditions of non-degeneracy, there exists a local neighborhood around the fixed point where the priority policy set is affine. This is the key observation which leads to the exponentially small error bounds under simple local stability conditions. This observation around the fixed point drove several works (Hong et al., 2023, 2024a, 2024b; Goldsztajn and Avrachenkov, 2024; Yan, 2024) to shift their perspective towards policies that look to steer the dynamical system towards the optimal fixed point. Hong et al. (2023) created a virtual system that was driven to the fixed point. To transform this into an asymptotically optimal policy, they required the real system to synchronize with the virtual system, hence, they developed the synchronization assumption. Unlike Hong et al. (2023) which passively allowed arms to align with each other eventually, Hong et al. (2024a) used a two set policy to actively align non aligned arms to the fixed point. In contrast to this form of control, Yan (2024) designed the align and steer policy to steer the mean field dynamical control system towards the fixed point. Yan (2024)’s policy is asymptotically optimal under the assumptions of controllability of the dynamical system to more directly move the state of the system towards the fixed point over time. In parallel to this work, Goldsztajn and Avrachenkov (2024) used a similar idea of mean field dynamical control for weakly coupled systems under a relaxed but generalized constraint set if the single arm process is unichain and aperiodic.
@@
From an algorithmic perspective, there are also a lot of recent papers on RMABs for the finite-horizon reward criteria Hu and Frazier (2017); Zayas-Caban et al. (2019); Brown and Smith (2020); Ghosh et al. (2022); Zhang and Frazier (2021); Gast et al. (2023a), or the infinite-horizon discounted reward Zhang and Frazier (2022); Ghosh et al. (2022). Our results are close in spirit to those since our policy shows how to use a finite-horizon policy to construct a policy that is asymptotically optimal for the infinite-horizon rewards.
Finally, it is worth noting that a recent result Yan et al. (2024), developed in parallel to our own showed an order $\mathcal{O}(1/N)$ result for the degenerate finite time horizon problem using a diffusion approximation instead of a mean field approximation. Unfortunately, this finite horizon result cannot immediately be compared to the infinite horizon average reward result primarily because unlike the mean field solution which can directly be translated as the solution to (5) the diffusion dynamic has no immediate corresponding equivalent.
@@ B.1 Rounding procedure
    First, given $\bm{u}=\mu_{\tau}({\bm{x}}(\textbf{S}(t)))$, we need to construct a vector $\bm{U}^{N}$ that is as close as possible to $\bm{u}$ while satisfying the constraints that $\sum_{i}u_{i}\leq\alpha$ and $N\bm{U}^{N}$ is an integer.
@@
This leads us to construct Algorithm 2. This algorithm first construct a $\bm{v}$ that is as close as possible to $\bm{u}$ while satisfying $\|\bm{v}\|\leq\alpha$. There might be multiple choices for this first step and the algorithm can choose any. For instance, applied to the example such that ${\bm{x}}=(10,10,10,9)$ and $\bm{u}=(10,4.9,4.6,0)$, this algorithm would first construct a vector $\bm{v}$ that can be any convex combination of $(10,4.4,4.6,0,0)$ and $(10,4.9,4.1,0,0)$. For instance, if the algorithm chose the vector $\bm{v}=(10,4.4,4.6,0)$, then the algorithm would produce $(10,5,4,0)$ with probability $0.4$ and $(10,4,5,0)$ with probability $0.6$. Once this is done, the algorithm outputs a $\bm{U}^{N}$ that satisfies the constraints and such that $\mathbb{E}[\bm{U}^{N}]=\bm{v}$. In some cases, there might be multiple distributions that work. The algorithm may output any. An efficient procedure to implemented this last step is provided in Section 5.2.3 of Ioannidis and Yeh (2016).
@@ C.1 Part 1, Properties of the dynamical control problem 8:
Proof C.3.
Now note, $\textbf{y}\cdot\textbf{1}=1$, where 1 is the all $1^{\prime}$s vector. We will overload the notation slightly by letting $\textbf{y}(t)$ be the optimal trajectory taken by the optimal policy $\pi^{*}_{\beta}$ for 19. Hence, for any constant $g$ we have:
@@
More importantly, by plugging these definitions into (30) and making the dependence of y on the maximal policy explicit, we obtain the following recursion equation:
@@
which completes the proof.
@@ C.2 Part 2: Dissipativity
which implies the results of the lemma.
@@ C.3 Part 3: MPC is optimal for the deterministic problem
Given any initial distribution ${\bm{x}}\in\Delta_{\mathcal{S}}$, the cost minimization problem $L_{\tau}({\bm{x}})$ and the reward maximization problem $W_{\tau}{({\bm{x}})}$ are equivalent. Next, for all ${\bm{x}}$, the limit $\lim_{t\to\infty}L_{t}({\bm{x}})$ exists and is bounded. Further, for any $\epsilon>0$, there exists $\tau(\epsilon)<\infty$ such that:
$|\lim_{t\to\infty}L_{t}({\bm{x}})-L_{\tau(\epsilon)}({\bm{x}})|<\epsilon.$
Finally, the limit so defined is Lipschitz with Lipschitz constant $C_{L}$ bounded above by $\|\lambda\|_{\infty}+\frac{k}{\rho_{k}}$.
Proof C.7.
@@
[…]
By Lemma  C.4, the rotated cost is $\tilde{l}(\textbf{x},\bm{u})=g^{*}-R({\bm{x}},\bm{u})+\lambda({\bm{x}})-\lambda(\Phi({\bm{x}},\bm{u}))$. As any valid control of $L_{\tau}$ and $W_{\tau}$ satisfy ${\bm{x}}(t+1)=\Phi({\bm{x}}(t),\bm{u}(t))$, the objective of $L_{\tau}$ can be rewritten as:
@@
As the last two terms correspond to the objective function for $W_{\tau}$, this shows that
$L_{\tau}({\bm{x}})=\tau g^{\star}+\lambda\cdot{\bm{x}}-W_{\tau}{({\bm{x}})}.$ (36)
Clearly, the two objectives are equivalent and $\mu_{\tau}({\bm{x}})=\bm{u}(0)$. Now note, by Lemma  C.2 we have,
@@
is bounded. Due to dissipativity, $L_{\tau}(\textbf{x})$ is monotone increasing in $\tau$. Recall, we denote the limit as $\tau\to\infty$ of the cost minimization problem by $L_{\infty}({\bm{x}})$. It follows that there exists a $\tau(\epsilon)$ such that,
$|L_{\infty}({\bm{x}})-L_{\tau(\epsilon)}({\bm{x}})|<\epsilon$
. Finally, note that:
$L_{\infty}({\bm{x}})=\lambda\cdot{\bm{x}}+h^{\star}({\bm{x}}).$
Hence, the Lipschitz constant under the $l1$ norm can explicitly be bounded as follows,
$|L_{\infty}({\bm{x}})-L_{\infty}({\bm{x}}^{\prime})|$ $\leq|\lambda\cdot({\bm{x}}-{\bm{x}}^{\prime})|+|h^{\star}({\bm{x}})-h^{\star}({\bm{x}}^{\prime})|$
$\leq\|\lambda\|_{\infty}\|{\bm{x}}-{\bm{x}}^{\prime}\|_{1}+\frac{k}{\rho_{k}}\|{\bm{x}}-{\bm{x}}^{\prime}\|_{1}$
$=\left(\|\lambda\|_{\infty}+\frac{k}{\rho_{k}}\right)\|{\bm{x}}-{\bm{x}}^{\prime}\|_{1}$
Where we used Holder’s inequality and Lemma C.2 for the second inequality.
C.4 Proof of Theorem 4.1: Computation details for (17)
@@
$\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[L_{\infty}({\bm{X}(t)})-L_{\infty}(\Phi({\bm{X}(t)},\bm{u}(t)))+2\epsilon-\lambda\cdot{\bm{X}}(t)+\lambda\cdot\Phi({\bm{X}}(t),\bm{u}(t))\right],$ (37)
where we use the definition of the rotated cost for the first equality, the dynamic principle for the second line and Lemma C.6 for the last line by choosing $\tau=\tau(\epsilon)$.
$\frac{1}{T}\sum_{t=0}^{T-1}[L_{\infty}({\bm{X}(t)})-\lambda\cdot{\bm{X}(t)}]$ $=\frac{1}{T}\sum_{t=0}^{T-1}[L_{\infty}({\bm{X}(t+1)})-\lambda\cdot{\bm{X}(t+1)}]$
$\quad+\frac{1}{T}\left(L_{\infty}({\bm{X}(0)})-L_{\infty}({\bm{X}(T)})+\lambda\cdot{\bm{X}(0)}-\lambda\cdot{\bm{X}(T)}\right)$
@@
$2\epsilon+\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[L_{\infty}({\bm{X}}(t+1))-L_{\infty}(\Phi({\bm{X}}(t),\bm{u}(t)))-\lambda\cdot{\bm{X}}(t+1)+\lambda\cdot\Phi({\bm{X}}(t),\bm{u}(t))\right],$
which is (17)
C.5 Proof of Theorem 4.1: Final computation details
In Lemma C.6, we show that the function $L_{\infty}(\cdot)$ has the following property in the $l1$ norm: $|L_{\infty}({\bm{x}})-L_{\infty}({\bm{x}}^{\prime})|\leq(\|\lambda\|_{\infty}+k/\rho_{k})\|{\bm{x}}-{\bm{x}}^{\prime}\|_{1}$. Moreover, we show that under appropriate parametrization (see Appendix F), the Lagrange multiplier $\|\lambda\|_{\infty}$ is bounded above by $\frac{k}{\rho_{k}}\left(1+\frac{\alpha k}{\rho_{k}}\right)$. This shows that
$|\mathbb{E}\Big{[}L_{\infty}({\bm{X}}(t$ $+1))-L_{\infty}(\Phi({\bm{X}}(t),\bm{u}(t)))-\lambda\cdot{\bm{X}}(t+1)+\lambda\cdot\Phi({\bm{X}}(t),\bm{u}(t))\Big{]}|$
$\leq(k/\rho_{k}+2\|\lambda\|_{\infty})\mathbb{E}\left[\|{\bm{X}}(t+1)-\Phi({\bm{X}}(t),\bm{u}(t)\|_{1}\right]$
$\leq\left(\frac{k}{\rho_{k}}+2\frac{k}{\rho_{k}}\left(1+\frac{\alpha k}{\rho_{k}}\right)\right)\mathbb{E}\left[\|{\bm{X}}(t+1)-\Phi({\bm{X}}(t),\bm{u}(t))\|_{1}\right]$
$\leq\frac{k}{\rho_{k}}\left(3+\frac{2\alpha k}{\rho_{k}}\right)\mathbb{E}\left[\|{\bm{X}}(t+1)-\Phi({\bm{X}}(t),\bm{u}(t))\|_{1}\right]$
To study the above equation, by the triangular inequality, we have that:
$\mathbb{E}\left[\|{\bm{X}}(t+1)-\Phi({\bm{X}}(t),\bm{u}(t))\|_{1}\right]$
$\leq\mathbb{E}\left[\|{\bm{X}}(t+1)-\Phi({\bm{X}}(t),\bm{U}(t))\|_{1}\right]+\mathbb{E}\left[\|\Phi({\bm{X}}(t),\bm{U}(t)-\Phi({\bm{X}}(t),\bm{u}(t))\|_{1}\right].$
[…]
$\mathbb{E}\left[\|{\bm{X}}(t+1)-\Phi({\bm{X}}(t),\bm{u}(t)\|_{1}\right]\leq\frac{\sqrt{|S|}}{\sqrt{N}}+2\frac{\alpha N-\lfloor\alpha N\rfloor}{N}.$
Plugging this into (17) shows that
$(A)\leq 2\epsilon+\frac{k}{\rho_{k}}\left(3+2\frac{\alpha k}{\rho_{k}}\right)\left(\frac{\sqrt{|S|}}{\sqrt{N}}+2\frac{\alpha N-\lfloor\alpha N\rfloor}{N}\right).$
By (16), $V_{\text{opt}}^{\textbf{x}}{N}-V_{\text{LP -}\tau}^{(N)}(\textbf{x})\leq(A)+\frac{\alpha N-\lfloor\alpha N\rfloor}{N}$. Hence:
$V_{\text{opt}}^{\textbf{x}}{N}-V_{\text{LP -}\tau}^{(N)}(\textbf{x})\leq 2\epsilon+\frac{k}{\rho_{k}}\left(3+2\frac{\alpha k}{\rho_{k}}\right)\left(\frac{\sqrt{|\mathcal{S}|}}{\sqrt{N}}+\frac{2(\alpha N-\lfloor{\alpha N}\rfloor)}{N}\right)+\frac{(\alpha N-\lfloor{\alpha N}\rfloor)}{N}$
Theorem 4.1 holds by grouping the terms that depend on $\frac{(\alpha N-\lfloor{\alpha N}\rfloor)}{N}$ and the terms that depend on $1/\sqrt{N}$.
@@
The lemma can then be proven by adapting the proof of Theorem 3 of Gast et al. (2023b) and in particular their Equation (17). This is proven by using Lemma 1 Gast et al. (2023b) that implies that $\Pr[\|{\bm{X}(t+1)}-\Phi({\bm{X}(t)},\bm{U}(t))\|\geq\epsilon]\leq e^{-C^{\prime}N}$.
We are now ready to prove Theorem 4.2. We can follow the proof of Theorem 4.1 up to (17) that shows that $V_{\text{opt}}^{\textbf{x}}{N}-V_{\text{LP -}\tau}^{(N)}(\textbf{x})$ is bounded [^5] In this expression, the term $(\alpha N-\lfloor{\alpha N}\rfloor)/N$ is equal to 00 here because we assumed that $\alpha N$ is an integer. by
@@
Proof Outline for Lemma C.1
The first part of this section is dedicated to proving the Lipschitz property for the value function $V^{\infty}({\bm{x}})$ in ${\bm{x}}$ under the ergodicity Assumption 1. For simplicity of exposition we will set $k=1$, although it is not hard to extrapolate the proof for $k>1$ and is left as an exercise for the reader.